# generalised polynomial chaos

### v2

@@ I Introduction
distributed energy resources are increasingly participating in ancillary service (AS) markets. Many of these resources. such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities. To facilitate the participation of these DERs, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [1]. This has important consequences for the reliable procurement of reserve capacity. This raises new questions with respect to the amount of reserve that should be procured, as the availability of the procured reserve becomes a stochastic quantity.
@@
Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables. Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions. This approach was first used for stochastic power flow computations [5], with inequality constraints formulated only with respect to the expected value. Later work, such as [6], introduces the use of gPC for the solution of optimal power flow problems. Typically, approaches modelling uncertainty with gPC formulate chance constraints through a moment-based reformulation. For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [7]. For other distributions, [3] proposes an iterative tuning approach for the bounds in the moment-based reformulation in an attempt to recover the desired constraint satisfaction rate. A major limitation of these approaches, is that they consider the different random input variables to be independent. To allow modelling dependencies between random variables, [8] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure. A challenge with this approach is that it leads to numerical issues due to the poor conditioning of the GS procedure[2]. Other methods to model dependent random variables with gPC exist, as reported for example in [2]. The Rosenblatt transformation is a mapping that transforms dependent random variables to a space of independent random variables, in which an orthonormal polynomial basis can be constructed using standard approaches. This has for example been applied to solve stochastic power flows [11]. However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients. For linear dependencies, this can be simplified to the Nataf transform [4], but the authors of [2] show that it performs worse than their proposed GS procedure. Other approaches, so-called domination methods, approximate functions of dependent random variables with functions of independent random variables, but only show good performance when an accurate approximation can be obtained [2].
@@
None of the works above handling reserve procurement allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies. Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time. Therefore, this paper makes the following contributions:
@@
    A convex reformulation for probabilistic power reserve procurement considering dependent random variables.
@@ II Methodology
Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{\top}$ of dimension $d$ with joint probability density function (PDF) $p(\bm{\xi})$ and support $\Xi$. As shown in previous works (e.g. [8]), a square-integrable random function $q(\bm{\xi})$ can be represented by a PCE
@@
where the equality becomes exact as the arbitrary number of terms $L$ goes to infinity. The basis polynomials $\{\psi_{l}\}$ are orthonormal with respect to the inner product
$\left\langle\psi_{m},\psi_{n}\right\rangle=\int_{\Xi}\psi_{m}(\bm{\xi})\,\psi_{n}(\bm{\xi})\,p(\bm{\xi})\,\mathrm{d}\bm{\xi}.$ (2)
The expansion coefficients $\{a_{l}\}$ are obtained by projection onto the polynomial basis. Since this inner product corresponds to an expectation, the coefficients read
$a_{l}=\left\langle q,\psi_{l}\right\rangle=\mathbb{E}\!\left[q(\bm{\xi})\,\psi_{l}(\bm{\xi})\right].$ (3)
The main challenge lies in the construction of orthonormal polynomial bases with respect to an arbitrary joint PDF $p(\bm{\xi})$. For independent random variables, orthogonal polynomial families associated with the Askey scheme can be employed, since they are orthogonal with respect to their corresponding marginal distributions (e.g. Hermite polynomials for Gaussian random variables).
For general dependent random variables, the basis can be constructed from a suitably chosen set of multivariate monomials of the random variables (e.g. all monomials with degree $\leq\nu$). As noted, a GS procedure can be used for the construction, but is numerically ill-conditioned [2]. To address this, [10] provides an alternative algorithm based on a Whitening transformation, relying only in part on GS procedures.
Finally, the moments of the output distribution of $q(\bm{\xi})$ can be recovered from the expansion coefficients. For instance the mean and variance are given by
@@ II-B Joint Probability using Copulas
For practical modelling of statistically dependent random variables, the joint PDF $p(\bm{\xi})$ must be approximated in a tractable manner. This requires specifying both the marginal distributions of the individual components $f(\xi_{i})$ as well as their dependence structure.
@@
Sklar’s Theorem states, that for every multivariate CDF there exists a unique Copula capturing the general dependence structure given by (5). In practice, one of several parametric Copula families is used.
For the application considered in this paper no specific tail-dependence is expected. As a result, a Gaussian Copula is selected to maintain computational tractability and leverage its closed-form mathematical expressions. Identifying $u_{i}=F_{i}(\xi_{i})$, it is defined as:
@@
where $\Phi_{\bm{\Sigma}}(\bm{z})$ is the multivariate normal CDF with correlation matrix $\bm{\Sigma}$ and $\Phi^{-1}(u_{i})$ the inverse normal CDF. The joint PDF is then expressed as
@@
where $c(\bm{u})$ is the Copula density. For a Gaussian Copula, the density is given by
$c_{\bm{\Sigma}}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
where $\phi_{\bm{\Sigma}}(\bm{x})$ is the multivariate normal PDF with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
@@ II-C Quadrature Integration
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}$ (9)
with $g(\bm{\xi})$ an arbitrary function. In general, these integrals cannot be evaluated analytically and must be approximated numerically. Given the potentially high dimensionality of the integrals, Monte Carlo sampling is a natural choice. This however requires a large amount of samples to achieve sufficient accuracy and sampling errors may accumulate throughout the PCE process. Moreover, the use of importance sampling is not straightforward in this context, since the integrands encountered vary across polynomial orders and random functions.
@@
$u_{i}=F_{i}(\xi_{i})\in[0,1],\qquad i=1,\dots,d.$ (10)
The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$ the marginal PDFs. Hence
$d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}.$ (11)
@@
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}g\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
where $\bm{F}^{-1}(\bm{u})$ is short-hand notation for the componentwise inverse marginal CDF. Evidently, this transformation is valid no matter the Copula used, in principle allowing any Copula family to be used for this quadrature integration scheme. However, the Copula density $c(\bm{u})$ is in general neither polynomial nor constant, meaning Gauss-Legendre quadrature will be inexact.
For the specific case of Gaussian Copulas, consider a second transformation to a Gaussian latent space
$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d$ (13)
on which tensor-grid Gauss-Hermite quadrature is applicable. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
@@
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\bm{\Phi}(\bm{z}))\big)\;\phi_{\bm{\Sigma}}(\bm{z})\,d\bm{z}$ (15)
with $\bm{\Phi}(\bm{z})$ the componentwise normal CDF. The result in (15) features a correlated normal PDF. However, Gauss-Hermite quadrature expects to weighted by an uncorrelated normal PDF. To address this, let $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ be the Cholesky factorisation of the correlation matrix. For a given set of quadrature nodes $\{\bm{z}_{n}\}$ and weights $\{w_{n}\}$ the linear transformation
@@
allows to “colour“ the correlation onto the quadrature nodes resulting in
@@
The number of nodes is given by $N=k^{d_{I}}$ with $k$ the number of evaluation points per dimension and $d_{I}$ the integral dimensionality. Due to this exponential scaling, dimensionality reduction strategies are necessary. Note that while the second transformation eliminated the non-polynomial Copula density $c(\bm{u})$, improving numerical accuracy, there is still no exactness guarantee. Gauss-Hermite quadrature is exact for polynomials up to degree 2k12k-1 weighted by a normal PDF, but $g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz})))$ does not necessarily behave like a polynomial of degree $\leq 2k-1$ in Gaussian latent space.
@@ II-C1 Numerical Improvements
where $p_{1,2}(\xi_{1},\xi_{2})$ is a reduced joint PDF. For elliptic Copulas they have identical structure to the full joint PDF defined in (7), just with less input variables. For other Copulas, the reduced joint PDFS might need to be retrieved separately.
Now, instead of directly applying an orthonormalization scheme, the monomial expectation values of all pairs of monomials in the chosen set can be evaluated beforehand. When applying the orthonormalization scheme, the inner products can be expanded as in (18) and the expectation values substituted.
Crucially, precomputing expectation values can be done in a parallel fashion and naturally limits the maximum integral dimension to be evaluated to $2\nu$. This because the highest dimensional expectation values to be evaluated are of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$. This approach decouples the integral dimension $d_{I}$ from the random variable dimension $d$.
In the case where multiple random functions $q_{i}(\bm{\xi_{i}})$ only depending on subsets $\bm{\xi_{i}}$ of the random vector are expanded to the same polynomial basis, a similar approach can be used for the expansion coefficients. Here the expectation values $\mathbb{E}[q_{i}(\bm{\xi_{i}})],\ldots,\mathbb{E}[q_{i}(\bm{\xi_{i}})\xi_{d}^{\nu}]$ can be computed before with reduced dimension. Note, that the reduction in this case is dependent on the dimensionality of the random functions $q_{i}(\bm{\xi_{i}})$.
Finally, the set of monomials from which the polynomial basis is constructed can be tuned. For instance, assume the random functions $q_{i}(\bm{\xi_{i}})$ are models of PV power output at locations $i$ with $\bm{\xi_{i}}$ the local weather conditions like irradiation and temperature. Physically meaningless monomials such as the product of squares of the irradiations at two different locations can be dropped from the set without significantly affecting the expansion accuracy.
@@ III Use Case
Reserve power, procured from AS markets, is activated whenever the grid frequency exceeds or dips below certain thresholds to ensure stable frequency levels. Given the critical nature of reserve activation, the risk of under-delivery must be kept to a minimum. This need for security can lead to high procurement costs. A recent initiative to keep economic efficiency is the Allocation of Cross-zonal Capacity and Procurement of aFRR Cooperation Agreement (ALPACA) which co-optimizes the procurement of reserves across borders of Austria, Czechia and Germany.
To validate the formalism introduced in this paper, a simple cross-border joint procurement optimization with uncertain bids and chance constraints on the available power is solved.
Formally, let $X$ and $Y$ be two zones sharing a common market with a set of bids $E_{i}(\bm{\xi})$ and costs $c_{i}$. Bids $i={1,\ldots,N_{X}}$ are located in zone $X$ and $i={N_{X}+1,\ldots,N_{X}+N_{Y}}$ are located zone Y. The decision variables $x_{i},y_{i}\in[0,1]$ determine the procured share of each bid. The objective is to minimize total cost, while both zones meet their reserve needs $R^{(X)},R^{(Y)}$ and stay below exchange limits $T_{X\to Y},T_{Y\to X}$ with $99\%$ probability. Expressed in standard form
$\min_{\bm{x},\bm{y}}$ $\bm{c^{T}}(\bm{x}+\bm{y})$ (20)
@@
$\sum_{j=1}^{Z}z_{j}E_{j}(\bm{\xi})$ $=\sum_{j=1}^{Z}z_{j}\sum_{l=0}^{L}a_{l}^{j}\psi_{l}(\bm{\xi})$ (21)
$=\sum_{l=0}^{L}\bar{a}_{l}\psi_{l}(\bm{\xi})$
@@
where $\lambda_{1}$ is the quantile factor for which $\mu+\lambda_{1}*\sigma$ corresponds to the $1^{st}$ percentile of the distribution. For normal distributions $\lambda_{1},\lambda_{99}=\mp 2.326$. Applying this reformulation to all chance-constraints in problem (20) results in a convex optimization problem.
@@
We solved a concrete example with $N_{X},N_{Y}=4$ bids per zone and $d=8$ random variables. In two situations. The bid functions are assumed to represent PV and HP capabilities. The random variables stand for the temperature and irradiation at four different locations with each two bids. Fig. 1 illustrates the setup.

### v3

@@ I Introduction
DER are increasingly participating in ancillary services markets. Many of these resources, such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities. To facilitate the participation of these distributed energy resources, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [4]. This raises new questions with respect to the amount of ASs that should be procured, as the full availability of the contracted ASs becomes uncertain. Existing works considering probabilistic reserve procurement either consider distributionally robust methods to account for uncertainties (e.g. [15] or [13]) or formulate chance constraints for the availability of different reserve bids, but do not account for the dependencies between the stochastic bid availabilities (e.g. [5]).
@@
Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model the propagation of uncertain input variables. Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions. This approach was first used for stochastic power flow computations [9], with inequality constraints formulated only with respect to the expected value. Later work, such as [10], introduces the use of gPC for the solution of optimal power flow problems. Typically, approaches modelling uncertainty with gPC formulate chance constraints through a moment-based reformulation. For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [11]. For other distributions, iterative tuning approaches for the bounds in the moment-based reformulation exist [7] attempting to recover the desired constraint satisfaction rate. A major limitation of these approaches is, that they consider the different random input variables to be independent. This can lead to underestimating the probability of violating certain thresholds. To allow modelling dependencies between random variables, [12] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure. A challenge with this approach is that it leads to numerical issues due to the poor conditioning of the GS procedure[6]. Other methods to model dependent random variables with gPC exist, as reported for example in [6]. The Rosenblatt transformation is a mapping that transforms dependent random variables to a space of independent random variables, in which an orthonormal polynomial basis can be constructed using standard approaches. This has for example been applied to solve stochastic power flows [16]. However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients. For linear dependencies, this can be simplified to the Nataf transform [8], but the authors of [6] show that it performs worse than their proposed GS procedure. Other approaches, so-called domination methods, approximate functions of dependent random variables with functions of independent random variables, but only show good performance when an accurate approximation of the orthogonal basis in terms of inner products with respect to the original joint distribution can be obtained [6].
@@
None of the works above handling reserve procurement allow to consider uncertain bids exhibiting dependencies. Additionally, the proposed approaches to extend gPC to dependent variables lead to computational issues due to error propagation and computation time. Therefore, this paper makes the following contributions:
@@
    A convex reformulation for probabilistic AS procurement considering dependent random variables.
@@ II Methodology
Consider the $d$-dimensional random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{\top}$ with joint probability density function (PDF) $p(\bm{\xi})$ and support $\Xi$. As shown in previous works (e.g. [12]), a square-integrable random function $q(\bm{\xi})$ can be represented by a PCE.
@@
with the equality becoming exact as the arbitrary number of terms $L$ goes to infinity. The basis polynomials $\{\psi_{l}\}$ are orthonormal with respect to the inner product
$\left\langle\psi_{m},\psi_{n}\right\rangle=\int_{\Xi}\psi_{m}(\bm{\xi})\,\psi_{n}(\bm{\xi})\,p(\bm{\xi})\,\mathrm{d}\bm{\xi}$ (2)
and the expansion coefficients $\{a_{l}\}$ are obtained by projection onto the polynomial basis. Since this inner product corresponds to an expectation, the coefficients read
$a_{l}=\left\langle q,\psi_{l}\right\rangle=\mathbb{E}\!\left[q(\bm{\xi})\,\psi_{l}(\bm{\xi})\right]$ (3)
The moments of the output distribution of $q(\bm{\xi})$ can be recovered from the expansion coefficients. For instance, the mean $\mu$ and variance $\sigma$ are given by
@@
The main challenge lies in the construction of orthonormal polynomial bases with respect to an arbitrary joint PDF $p(\bm{\xi})$. For independent random variables, polynomial families associated with the Askey scheme can be employed, since they are orthogonal with respect to their corresponding marginal distributions (e.g. Hermite polynomials for Gaussian random variables).
For general dependent random variables, the basis can be constructed from a suitably chosen set of multivariate monomials of the random variables (e.g. all monomials with degree $\leq\nu$). As noted in the introduction, a GS procedure can be used for the construction, but is numerically ill-conditioned [6]. To address this, [14] provides an alternative algorithm based on a Whitening transformation, relying only in part on GS procedures.
@@ II-B Joint Probability using Copulas
For practical modelling of statistically dependent random variables, the joint PDF $p(\bm{\xi})$ must be approximated in a tractable manner. This requires specifying both the marginal distributions of the individual components $f_{i}(\xi_{i})$ as well as their dependence structure.
@@
Sklar’s Theorem states that for every multivariate CDF there exists a Copula capturing the general dependence structure given by (5). In practice, one of the several parametric Copula families is used.
For the application considered in this paper, a Gaussian Copula is selected. The Gaussian Copula captures dependence through a correlation matrix $\bm{\Sigma}$ containing the pairwise correlations between the random variables. Identifying $u_{i}=F_{i}(\xi_{i})$, it is defined as:
@@
where $\Phi_{\bm{\Sigma}}(\bm{z})$ is the multivariate normal CDF with correlation matrix $\bm{\Sigma}$ and $\Phi^{-1}(u_{i})$ the inverse normal CDF. This closed-form expression allows efficient scaling to higher dimensions while maintaining computational tractability. In contrast, Archimedean Copulas provide limited flexibility in higher dimensions due to their single-parameter structure, whereas Vine Copulas can represent complex and asymmetric dependence structures but rapidly become intractable as dimensionality increases. Using a Copula, with its density expressed as $c(\bm{u})$, the joint PDF can be formulated as
@@
For a Gaussian Copula, the density is given by (8), where $\phi_{\bm{\Sigma}}(\bm{x})$ is the multivariate normal PDF with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
$c_{\bm{\Sigma}}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)}$ (8)
@@ II-C Quadrature Integration
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})d\bm{\xi}$ (9)
with $g(\bm{\xi})$ an arbitrary function. In general, these integrals cannot be evaluated analytically and must be approximated numerically. Given the potentially high dimensionality of the integrals, Monte Carlo sampling is a natural choice. This, however, requires a large amount of samples to achieve sufficient accuracy and sampling errors may accumulate throughout the PCE process. Moreover, the use of importance sampling is not straightforward in this context, since the integrands encountered vary across polynomial orders and random functions.
@@
$u_{i}=F_{i}(\xi_{i})\in[0,1]\qquad i=1,\dots,d$ (10)
The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$ the marginal PDFs. Hence,
$d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}$ (11)
@@
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}g\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
where $\bm{F}^{-1}(\bm{u})$ is short-hand notation for the component-wise inverse marginal CDF. Evidently, this transformation does not rely on the specific Copula used, in principle allowing any Copula family to be used for this quadrature integration scheme. However, the Copula density $c(\bm{u})$ is in general neither polynomial nor constant, meaning Gauss-Legendre quadrature will be inexact. For the specific case of Gaussian Copulas, consider a second transformation to a Gaussian latent space
$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty)\qquad i=1,\dots,d$ (13)
on which tensor-grid Gauss-Hermite quadrature is applicable. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, meaning
@@
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\bm{\Phi}(\bm{z}))\big)\;\phi_{\bm{\Sigma}}(\bm{z})\,d\bm{z}$ (15)
with $\bm{\Phi}(\bm{z})$ the component-wise normal CDF. The result in (15) features a correlated normal PDF. However, Gauss-Hermite quadrature expects to weighted by an uncorrelated normal PDF. To address this, let $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ be the Cholesky factorisation of the correlation matrix. For a given set of quadrature nodes $\{\bm{z}_{n}\}$ and weights $\{w_{n}\}$ the linear transformation
@@
allows to “color“ the correlation onto the quadrature nodes resulting in
@@
The number of nodes is given by $N=k^{d_{I}}$ with $k$ the number of evaluation points per dimension and $d_{I}$ the integral dimensionality. Due to this exponential scaling, dimensionality reduction strategies are necessary. Note that, even though the second transformation eliminates the non-polynomial Copula density $c(\bm{u})$ and improves numerical accuracy, there is still no exactness guarantee. Tensor-grid Gauss–Hermite quadrature is exact only for integrands that are polynomials up to degree 2k12k-1 in each variable weighted by a normal PDF, but $g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz})))$ does not necessarily behave like a polynomial in Gaussian latent space. For a comprehensive treatment of quadrature rules and their exactness properties, see classical texts such as [1].
@@ II-C1 Numerical Improvements
where $\Xi_{1,2}$ is the corresponding support and $p_{1,2}(\xi_{1},\xi_{2})$ the joint PDF of the appearing random variables. For elliptic Copulas, like the Gaussian Copula, such lower dimensional joint PDFs have identical structure to the full joint PDF defined in (7) and are therefore easy to recover. For other Copulas, they might need to be computed separately.
Now, instead of directly applying an orthonormalization scheme, the expectation values of all pairwise products of monomials in the chosen set can be evaluated beforehand. When applying the orthonormalization scheme, the inner products can be expanded as in (18) and the expectation values substituted. Crucially, precomputing expectation values can be done in a parallel fashion and naturally limits the maximum integral dimension to be evaluated to $2\nu$, as the highest dimensional expectation values to be evaluated are of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$. This decouples the integral dimensionality $d_{I}$ from the random variable dimension $d$. In the case where multiple random functions $q_{i}(\bm{\xi_{i}})$ only depending on subsets $\bm{\xi_{i}}$ of the random vector are expanded to the same polynomial basis, a similar approach can be used for the expansion coefficients. Here the expectation values $\mathbb{E}[q_{i}(\bm{\xi_{i}})],\ldots,\mathbb{E}[q_{i}(\bm{\xi_{i}})\xi_{d}^{\nu}]$ can be computed beforehand with reduced dimension. Note, that the reduction in this case is dependent on the dimensionality of the random functions $q_{i}(\bm{\xi_{i}})$. The implementation of the quadrature integration procedure using Gaussian Copulas is summarized in Algorithm 1.
Finally, the set of monomials from which the polynomial basis is constructed can be reduced to limit the computational cost. For instance, assume the random functions $q_{i}(\bm{\xi_{i}})$ are models of PV power output at locations $i$ with $\bm{\xi_{i}}$ the local weather conditions like irradiation and temperature. Physically irrelevant monomials, such as higher-order products of the irradiations at different locations, can be dropped from the set without significantly affecting the PCE accuracy. This approach is conceptually similar to sparse PCE, which discards basis polynomials with negligible influence on the expansion output [2].
@@ III Use Case
Reserve power, procured from AS markets, is activated whenever the grid frequency exceeds or dips below certain thresholds to ensure stable frequency levels. Given the critical nature of reserve activation, the risk of under-delivery must be kept to a minimum. This need for security can lead to high procurement costs. A recent initiative to keep economic efficiency is the Allocation of Cross-zonal Capacity and Procurement of aFRR Cooperation Agreement (ALPACA) which co-optimizes the procurement of reserves across borders of Austria, Czechia and Germany. To validate the formalism introduced in this paper, a simple cross-border joint procurement optimization with uncertain bids and chance constraints on the available power is solved.
Formally, let $X$ and $Y$ be two zones sharing a common AS market with a set of bids $E_{i}(\bm{\xi})$ and costs $\gamma_{i}$. Bids $i={1,\ldots,N_{X}}$ are located in zone $X$ and $i={N_{X}+1,\ldots,N_{X}+N_{Y}}$ are located zone Y. The decision variables $x_{i},y_{i}\in[0,1]$ determine the procured share of each bid. The objective is to minimize total cost, while both zones meet their reserve needs $R^{(X)},R^{(Y)}$ and stay below exchange limits of the interconnecting tie lines $T_{X\to Y},T_{Y\to X}$ with $99\%$ probability. The value of this probability threshold is inspired from Art 157. (h)-(i) in [3]. Expressing the problem in standard form
$\min_{\bm{x},\bm{y}}$ $\bm{\gamma}^{\top}(\bm{x}+\bm{y})$ (20)
@@
Algorithm 1 Quadrature Integration with Dimension Reduction for a Gaussian Copula
0: Integrand function $g(\bm{\xi})$, correlation matrix $\bm{\Sigma}\in\mathbb{R}^{d\times d}$, number of evaluation points per dimension $k$
0: Numerical integral value II
1: Determine the effective dimension $d_{I}$ of $g(\bm{\xi})$,
2: Extract the submatrix $\bm{\Sigma}_{d_{I}}\in\mathbb{R}^{d_{I}\times d_{I}}$ corresponding to the relevant dimensions
3: Compute the Cholesky factorization $\bm{LL}^{\top}=\bm{\Sigma}_{d_{I}}$
4: Determine the $N=k^{d_{I}}$ tensor-grid Gauss-Hermite quadrature nodes $\{\bm{z_{n}}\}$ and weights $\{w_{n}\}$
[…]
6:for $n=1$ to NN do
7:  $I\leftarrow I+w_{n}\cdot g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz_{n}})))$
8:end for
9:
10:return II
@@
where $\lambda_{1}$ is the quantile factor for which $\mu+\lambda_{1}*\sigma$ corresponds to the $1^{st}$ percentile of the distribution. For normal distributions $\lambda_{1},\lambda_{99}=\mp 2.326$. Applying this reformulation to all chance-constraints in problem (20) results in the convex optimization problem (24).
@@
We consider a concrete example with $N_{X},N_{Y}=4$ bids per zone and $d=8$ random variables and analyze two distinct cases. The bids are distributed across four locations, with two bids at each location depending on two random variables that represent local conditions. Fig. 1 illustrates the use case.